how many liters of a 10 percent alcohol solution must be mixed with 70 liters of the solution that is 80 percent alcohol to get a solution that is 20 percent alcohol
Answer by mananth(16949)

percent ---------------- quantity
Alcohol 80 ---------------- 70
Alcohol 10 ---------------- x
Total 20 ---------------- 70+x
....
80*70+10 *x=20(70 +x)
5600+10x =1400+20 x
10x-20x= 1400-5600
-10x=-4200
/-10
x=420 liters 10% Alcohol